# PoSh: Using Scene Graphs to Guide LLMs-as-a-Judge for Detailed Image Descriptions

### v1

@@ A.3 Evaluation
A.3.1 Experimental Setting
Numbers for PoSh reported in Tables 2, 5 and 6 were produced on a single H100 GPU. Our codebase specifies the necessary versions of CUDA and dependent libraries like vLLM for replication (relying on vLLM’s implementation of batch invariance).
A.3.2 Correlation Metrics
@@
A.3.3 Coarse Metric Evaluation
@@ A.3.4 Granular Baselines
We extract all of the $4$-grams from each sentence of a generation and its reference, embed them using Qwen/Qwen3-Embedding-8B (Reimers & Gurevych, 2019; Yang et al., 2025) and then calculate the maximum pairwise similarities between generation $4$-grams and reference $4$-grams. Generation text spans and reference text spans with maximum pairwise similarity scores lower than $0.7$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for 4GramEmbed in Table 2.
@@
We extract all of the components (objects, attribute-object pairs, and object-relation-object triples) from the scene graphs of a generation and its reference extracted for PoSh in Section 3, embed them using Qwen/Qwen3-Embedding-8B (Reimers & Gurevych, 2019; Yang et al., 2025) and then calculate the maximum pairwise similarities between the generation components and the reference components. Generation components and reference components with maximum pairwise similarity scores lower than $0.8$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for SGEmbed in Table 2.
A.3.5 Coarse Baselines
@@ A.3.6 Reinforcement Learning
    DAPO (Yu et al., 2025) with full parameter updates, implemented with TRL (von Werra et al., 2020), using a learning rate of 1e61e-6, 2020 warmup steps, 88 generations per sample (with a temperature of $1.0$ and $top_{p}=0.7$), $\epsilon=0.28$, $\beta=0$, and an effective batch size of $64$ for a single epoch, choosing the final checkpoint
For both variants, we train using 8 H100 GPUs (Bloom et al., 2025).
@@ A.4.1 Coarse
Table 6: All coarse metrics evaluated on DOCENT and CapArena, identified with $\Theta$ (parameter count, in billions), [Uncaptioned image] (requires a reference), [Uncaptioned image] (requires an image) and [Uncaptioned image] (replicable). “acc” indicates accuracy at predicting the better generation (or “tie”) in each judged pair. For DOCENT, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between differences in the metric and differences in the rank of the generations in each pair. For CapArena, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between model ELO rankings derived from metric scores and human judgments. Bold indicates the best replicable metric while underlining indicates the best metric overall. Gray cells indicate correlations that are not statistically significant at $\alpha=0.05$. PoSh beats all replicable baselines and GPT4o on DOCENT in all settings (mistakes, omissions and overall quality) while remaining replicable. Moreover, PoSh is robust, achieving the second best score among replicable metrics on CapArena.
@@
length $\checkmark$ 30.5 -0.270 -0.206 37.8 -0.002 -0.001 38.0 -0.160 -0.121 58.7 0.710 0.582
BLEU-4 $\checkmark$ $\checkmark$ 34.2 -0.070 -0.053 42.5 0.118 0.087 42.8 0.051 0.038 47.4 0.424 0.319
CIDER $\checkmark$ $\checkmark$ 32.0 -0.118 -0.089 37.5 -0.009 -0.007 37.8 -0.106 -0.079 38.4 -0.279 -0.209
METEOR $\checkmark$ $\checkmark$ 36.0 -0.103 -0.078 46.2 0.260 0.197 44.8 0.113 0.084 57.6 0.785 0.582
@@
CLIPScore $\checkmark$ $\checkmark$ 45.3 0.145 0.108 47.0 0.176 0.133 53.5 0.181 0.136 32.5 -0.574 -0.451
FLEUR 13 $\checkmark$ $\checkmark$ 35.2 -0.053 -0.040 38.5 0.029 0.020 41.2 -0.040 -0.031 45.8 0.393 0.297
Prometheus 8x7 $\checkmark$ $\checkmark$ 51.2 0.014 0.011 49.8 0.136 0.116 58.5 -0.007 -0.007 53.9 0.859 0.648
@@
LLaVa Critic 72 $\checkmark$ $\checkmark$ $\checkmark$ 62.8 0.412 0.351 57.0 0.509 0.430 66.8 0.546 0.461 64.0 0.987 0.934
@@
GPT5 $\checkmark$ $\checkmark$ 68.2 0.604 0.494 56.3 0.477 0.366 67.2 0.602 0.475 62.1 0.934 0.846
PoSh 14 $\checkmark$ $\checkmark$ 60.3 0.518 0.405 61.8 0.578 0.449 70.3 0.600 0.466 59.2 0.921 0.780
Table 7: Annotator agreement (Krippendorff’s $\alpha$) and aggregate preferences between SFT and DAPO with PoSh. A PoSh generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. Reported numbers are averages over these scores.
@@
PoSh vs SFT 0.243-0.243 $0.432$ $0.135$

### v2

@@ A.3 Evaluation
A.3.1 Metrics
@@
A.3.2 Coarse Score Scaled Evaluation
@@ A.3.3 Granular Baselines
We extract all of the $4$-grams from each sentence of a generation and its reference, embed them using Qwen/Qwen3-Embedding-8B (reimers-2019-sentence-bert; yang2025qwen3) and then calculate the maximum pairwise similarities between generation $4$-grams and reference $4$-grams. Generation text spans and reference text spans with maximum pairwise similarity scores lower than $0.7$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for 4GramEmbed in Table 2.
@@
We extract all of the components (objects, attribute-object pairs, and object-relation-object triples) from the scene graphs of a generation and its reference extracted for PoSh in Section 3, embed them using Qwen/Qwen3-Embedding-8B (reimers-2019-sentence-bert; yang2025qwen3) and then calculate the maximum pairwise similarities between the generation components and the reference components. Generation components and reference components with maximum pairwise similarity scores lower than $0.8$ were predicted as mistakes and omissions respectively, a threshold chosen to maximize the macro F1 scores reported for SGEmbed in Table 2.
A.3.4 Coarse Baselines
@@ A.3.5 Reinforcement Learning
    DAPO (yu2025dapo) with full parameter updates, implemented with TRL (vonwerra2022trl), using a learning rate of 1e61e-6, 2020 warmup steps, $8$ generations per sample (with a temperature of $1.0$ and $top_{p}=0.7$), $\epsilon=0.28$, $\beta=0$, and an effective batch size of $64$ for a single epoch, choosing the final checkpoint
@@ A.4.1 Coarse
Table 6: All coarse metrics evaluated on DOCENT and CapArena, identified with $\Theta$ (parameter count, in billions), [Uncaptioned image] (requires a reference), [Uncaptioned image] (requires an image) and [Uncaptioned image] (replicable). “acc” indicates accuracy at predicting the better generation (or “tie”) in each judged pair. For DOCENT, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between differences in the metric and differences in the rank of the generations in each pair. For CapArena, $\rho$ / $\tau$ indicate the Spearman rank / Kendall’s $\tau$ correlations between model ELO rankings derived from metric scores and human judgments. Bold indicates the best replicable metric while underlining indicates the best metric overall. Gray cells indicate correlations that are not statistically significant at $\alpha=0.05$. PoSh beats all replicable baselines and GPT4o on DOCENT in all settings (mistakes, omissions and overall quality) while remaining perfectly replicable. Moreover, PoSh is robust, achieving the second best score among replicable metrics on CapArena.
@@
length $\checkmark$ $30.5$ 0.270-0.270 0.206-0.206 $37.8$ 0.002-0.002 0.001-0.001 $38.0$ 0.160-0.160 0.121-0.121 $58.7$ $0.710$ $0.582$
BLEU-4 $\checkmark$ $\checkmark$ $34.2$ 0.070-0.070 0.053-0.053 $42.5$ $0.118$ $0.087$ $42.8$ $0.051$ $0.038$ $47.4$ $0.424$ $0.319$
CIDER $\checkmark$ $\checkmark$ $32.0$ 0.118-0.118 0.089-0.089 $37.5$ 0.009-0.009 0.007-0.007 $37.8$ 0.106-0.106 0.079-0.079 $38.4$ 0.279-0.279 0.209-0.209
METEOR $\checkmark$ $\checkmark$ $36.0$ 0.103-0.103 0.078-0.078 $46.2$ $0.260$ $0.197$ $44.8$ $0.113$ $0.084$ $57.6$ $0.785$ $0.582$
@@
CLIPScore $\checkmark$ $\checkmark$ $45.3$ $0.145$ $0.108$ $47.0$ $0.176$ $0.133$ $53.5$ $0.181$ $0.136$ $32.5$ 0.574-0.574 0.451-0.451
FLEUR $13$ $\checkmark$ $\checkmark$ $35.2$ 0.053-0.053 0.040-0.040 $38.5$ $0.029$ $0.020$ $41.2$ 0.040-0.040 0.031-0.031 $45.8$ $0.393$ $0.297$
Prometheus 88x77 $\checkmark$ $\checkmark$ $51.2$ $0.014$ $0.011$ $49.8$ $0.136$ $0.116$ $58.5$ 0.007-0.007 0.007-0.007 $53.9$ $0.859$ $0.648$
@@
LLaVa Critic $72$ $\checkmark$ $\checkmark$ $\checkmark$ $\mathbf{62.8}$ $0.412$ $0.351$ $57.0$ $0.509$ $0.430$ $66.8$ $0.546$ $0.461$ $\mathbf{\underline{64.0}}$ $\mathbf{\underline{0.987}}$ $\mathbf{\underline{0.934}}$
DSG $\checkmark$ $41.3$ $0.091$ $0.068$ $37.3$ $0.033$ $0.024$ $45.3$ $0.017$ $0.012$ - - -
@@
GPT5 $\checkmark$ $\checkmark$ $\underline{68.2}$ $\underline{0.604}$ $\underline{0.494}$ $56.3$ $0.477$ $0.366$ $67.2$ $\underline{0.602}$ $\underline{0.475}$ $62.1$ $0.934$ $0.846$
PoSh $14$ $\checkmark$ $\checkmark$ $60.7$ 0.519 0.405 62.7 0.581 0.451 70.7 0.599 0.466 $59.2$ $0.931$ $0.796$
Table 7: Annotator agreement (Krippendorff’s $\alpha$) and aggregate preferences between variants of Qwen2.5-VL-7B trained on DOCENT: tuned with DAPO using PoSh (“PoSh-tuned”) and supervised fine-tuned (“SFT”). A PoSh-tuned generation earns a score between 2-2 and $2$ based on how much worse or better it is than its SFT counterpart. Reported numbers are averages of these scores.
@@
PoSh-tuned vs SFT 0.243-0.243 $0.432$ $0.135$
A.4.2 Subcomponent Performance
To evaluate the scene graph extraction subcomponent of PoSh, we hand annotate scene graphs for ten descriptions in DOCENT, five references and five generations. We measure precision, recall and F1 for entities, attributes and relations and, for each matched element pair (i.e. entity, attribute or relation), accuracies for coreference resolution, attribute attachment and relation head and tail attachment. The numbers reported in Table 8 are averaged across the ten descriptions.
To evaluate the scene graph element verification subcomponent of PoSh, we manually answer $620$ templated questions for two randomly sampled generation-reference pairs in DOCENT and compare them to PoSh’s presence scores. The numbers reported in Table 9 are the maximum achievable F1 scores across all alerting thresholds. The errors in this subcomponent stem from generations specifying correct details not present in the reference, relative permissiveness toward interpretive language in generations and limitations of the PoSh’s templating and unique identifier discovery logic. Nevertheless, the scores speak to the strength of the PoSh framework and the potential of future subcomponent improvements to yield further gains in performance.
Table 8: Evaluation of PoSh’s extracted scene graphs for ten descriptions in DOCENT. “P” indicates precision, “R” indicates recall, “coref” indicates accuracy at predicting each matched entity’s first mention, “head” indicates accuracy at predicting each matched attribute or relation’s entity subject and “tail” indicates accuracy at predicting each matched relation’s entity object.
Entities Attributes Relations
P R F1 coref P R F1 head P R F1 head tail
$0.937$ $0.883$ $0.909$ $0.925$ $0.881$ $0.917$ $0.898$ $0.951$ $0.978$ $0.794$ $0.872$ $0.850$ $0.926$
Table 9: Evaluation of PoSh’s scene graph verification for $620$ elements from two generation-reference pairs in DOCENT, broken out by mistakes (verifying generation scene graph elements in a reference) and omissions (verifying reference scene graph elements in a generation).
Mistakes F1 Omissions F1 Overall F1
$0.941$ $0.754$ $0.852$